OpenStudy (anonymous):

Answer This 2 Questions Get a Medal !!!!!!!!!

OpenStudy (anonymous):

How can translations help me understand how to represent verbal phrases as algebraic expressions? What are the key words and phrases that indicate certain operations?

jigglypuff314 (jigglypuff314):

It might have helped with a given example ;) so... for example given the vertex form of a parabola: y = (x - h)^2 + k when coordinate (h, k) is the vertex when it say something like "translate left 2 units" know that "left/right" would be along the x-axis therefore affecting the "h" variable in the vertex form. and when "up 3 units" know that "up/down" would be along the y-axis, therefore affecting the "k" variable in the vertex form. but be careful that when you do plug it into algebraic form that you take into account the negatives and such so given original : y = x^2 move to left 2 then it would become y = (x - (-2))^2 --> y = (x + 2)^2

OpenStudy (anonymous):

Hey I remember you you were with ummmm I forgot her name

jigglypuff314 (jigglypuff314):

LovelyAnna ;)

OpenStudy (anonymous):

yeah her sorry I didn't talk back I wen't somewhere with my mom after she when to get you

jigglypuff314 (jigglypuff314):

yeah that's fine ;) what would you like help with then?

OpenStudy (anonymous):

well this 2 questions right here it's difficult

jigglypuff314 (jigglypuff314):

well... translations affect the parent functions of algebraic expressions in such a way that using verbal phrases relates to how the function is translated. I think I covered some phrases in my first comment.

OpenStudy (anonymous):

wow your really smart what grade are you in?

OpenStudy (anonymous):

hey whats the other one

OpenStudy (anonymous):

or is that both answers

jigglypuff314 (jigglypuff314):

I'm failing math in 11th grade :P it should cover both answers... the first comment was for the second part

OpenStudy (anonymous):

ohh wow 11th grade math your almost graduating your lucky I want to get out of school and explore the life already not stuck in prison

OpenStudy (anonymous):

thanks by the way

jigglypuff314 (jigglypuff314):

you're welcome ;) and I suck at life, so I'll just stick with school thanks :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

well I'm going to do my homework if I need help I'll tell you ok :)

jigglypuff314 (jigglypuff314):

sure ;)